# Exploring Time Delay Interferometry Ranging as a Practical Ranging Approach in the Bayesian Framework

### v3

@@ II.1 TDIR on unsynchronized data
[…]
@@
[…]
@@
[…]
@@
which incorporates the information of both LTT and differential clock deviations between S/Cs. The first definition is expressed in TCB, and the second gives its transformation to the clock time of receiving S/C, as an application of Eq. (2). To effectively suppress LFN with TDI on unsynchronized data, similar to the “TDI without clock synchronization” scheme [32], the objective of TDIR can be interpreted as searching for the functional form that fits $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ best. However, perfect reconstruction of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ with a fitting function is not practical due to the presence of random clock jitters in $\delta\hat{\tau}_{i}^{t}$. Functional forms with limited number of parameters (as few as possible to ensure the efficiency of MCMC sampling) can at most fit the slowly-varying, deterministic part of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$. We refer to this variable as the “fiducial” TDIR $\bar{d}_{ij}^{\hat{\tau}_{i}}(\tau)$, to be distinguished it from the “parametrized” TDIR dubbed $R_{ij}(\tau,\bm{\theta})$. The fiducial TDIR will be treated as the “ground truth” of our searching algorithm, and can be calculated via Eq. (5) assuming that there is no random clock jitters. Meanwhile, the parametrized TDIR contains unknown model parameters to be fitted using the beatnote signals.
@@
[…]
@@
[…]
@@
For the applications of TDI, it is crucial that the delays used to shift the signals must be consistent with the delays appearing in the equations of signals. Formally Eq. (8-11) looks similar to their “LTT versions” [18, 34, 24, 25], thus all the deduction and algorithm of TDI still applies, and the only two alterations which should be made are to change the delays from LTT to TDIR, and to interpret the beatnote signals as the unsynchronized measurements. We therefore proved that TDIR can be used to suppress LFN on unsynchronized data.
By making slight modifications to the standard formulae in the literature, we explicitly present the expressions for the complete TDI data processing flow using TDIR. Firstly, the test-mass-to-test-mass interferometry (denoted as intermediate variable $\xi$) is constructed to eliminate the OB noise:
@@
[…]
@@ II.2 Requirements of noise suppression on TDIR
×$\times\left[P_{\rm OMS}(f)+P_{\rm ACC}(f)(\cos^{2}u+3)\right],$
@@
Next, we investigate the residual LRN, OB noise and clock noise caused by the ranging errors. For generality, we do not set a specific laser locking scheme [42, 38] but rather assume all the six lasers are independent. Besides, to derive analytical expressions for a given level of ranging accuracy, the errors of ranging are set as constant biases dubbed $\Delta R_{ij}$. According to Eq. (12-II.1), the PSDs of the residual LFN and OB noise in the $X_{2}$ channel are
@@
where $\omega\equiv 2\pi f$, $\nu_{0}=281.6\ {\rm THz}$ is the central frequency of laser, and
@@
defines the average ranging error. Eq. (18) is in consistency with the results of previous studies [32, 43], and one can easily verify that a $\mathcal{O}(10)\ {\rm ns}$ requirement is reasonable for the noise models in consideration. While, under the same $\Delta R$, the residual OB noise is much smaller. The algorithms of clock noise reduction varies in the literature (e.g. [34, 35, 39]), thus the residual clock noise can not be written in a unified analytical form. We adopt the treatment of Ref. [35], and explore the impact of ranging error via a clock-noise-only simulation.
The analytical and simulation results are visualized in FIG. 2. In our simulation, the ranging errors are generated according to normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises, which are 10 ns, 10 ns, 1 ms for the LFN, OB noise and clock noise, respectively. The grey, green and orange amplitude spectral density (ASD) plots represent the simulated residual LFN, OB noise and clock noise, with the corresponding theoretical results shown in black dashed curves. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression. Besides the consistency between theoretical analysis and simulations, the figure also indicates that although the ground-tracking ranges deviate from the fiducial TDIR for less than 1 ms, they are already sufficient for suppressing the clock noises. This result is beneficial for the practical implementation of TDIR. Currently the post-TDI clock noise algorithms are usually more complex than TDI itself, therefore require more computational time. To tackle this issue, we do not need to compute $X_{2,q}$ at every iteration of the MCMC run, but rather calculate it with the ground-tracking ranges in advance. Furthermore, within the regions of parameter space of our interest (tens of nanoseconds around the fiducial TDIR), OB and clock noises only make negligible contributions to the total noise budget, thus the performance of TDIR is mainly determined by LFN and the secondary noises.
@@ III TDIR in the Bayesian framework
Assuming that the instrumental noises are Gaussian and stationary, $\mathcal{L}$ can be expressed in the form of matched filtering likelihood:
@@
where $\tilde{X}(f;\bm{\theta})$ represents the Fourier transform of an arbitrary TDI channel named $X$, and $P_{X}(f;\bm{\theta})$ is its one-sided PSD (periodogram). The data is whitten by the theoretical one-sided PSD of the secondary noises $P^{\rm 2nd}_{X}(f)$, since secondary noises dominate the TDI data in the region of parameter space of our interest (tens of nanoseconds around the fiducial TDIR). In the realistic unequal-arm situation, $P^{\rm 2nd}_{X}(f)$ depends on the LTTs, which can take the values of ground-tracking ranges. It can be easily verified that neither the sub-millisecond uncertainties nor the variations of LTTs during a day have considerable impact on $P^{\rm 2nd}_{X}(f)$, using Eq. (40-44) via numerical or analytical calculation. Eq. (24) is equivalent to the ones adopted by previous studies [24, 25]. In the more general “global fit” [46] scenario where GW signal exists, ${\rm ln}\mathcal{L}$ should be replaced by a more familiar form $-1/2(X-h|X-h)$, $h$ being the template of GW signal.
Simultaneously estimating the delay parameters and GW source parameters can be a challenging task. Additionally, Taiji will be able to detect the GWs from tens of millions of compact binaries in the Milky Way, and the majority of them cannot be resolved individually, resulting in a foreground confusion noise in the mHz band [47]. Precise modeling of this noise is only possible when all the bright GW signals have been correctly recognized and subtracted. Therefore, to reduce the encounters with GW signals and foreground, we set $f_{\rm min}=10\ {\rm mHz}$, above the most sensitive band of Taiji. This choice of $f_{\rm min}$ also allows us to accelerate the estimation process by incorporating a preliminary search phase, see Sec. IV. Ideally, under the condition that the computing capacity is sufficient and the algorithm has been optimized, one may discard this phase and raise $f_{\rm min}$ above 0.1 Hz to further minimize the impact of GWs.
@@
Figure 4: The ASDs of secondary noises and residual LFN in the $A_{2}$ and $T_{2}$ channels, with the armlengths calculated from the orbit data in the 1st day, and the ranging error are set to 10 ns. The shapes of $E_{2}$ channel are similar to those of $A_{2}$ hence not shown.
@@
and the total logarithmic likelihood function is the sum of $\{A_{2},E_{2},T_{2}\}$ channels. The PSDs of secondary noises in the optimal channels can be found in Appendix A, where we have also presented the PSDs of residual LFN due to the coupling with ranging errors. Observing Eq. (25), the contribution of a certain TDI channel to the total likelihood is determined by the ratio between residual LFN and secondary noises. These two terms for the $\{A_{2},T_{2}\}$ channels are shown in FIG. 4, with the armlengths calculated from the numerical orbit data in the 1st day, and the ranging errors are set to 10 ns. The shapes of noises in $E_{2}$ are similar to those of $A_{2}$ hence not shown. As can be seen from the figure, within the frequency band in concern, each channel contributes similarly to the likelihood.
[…]
@@ IV.2 Results of MCMC sampling
Figure 6: Left panel: The distributions of $\Delta R$ calculated from the MCMC samples of Simulation I, II. $\Delta R_{\rm median,I}=0.49\ {\rm ns}$ and $\Delta R_{\rm median,II}=5.28\ {\rm ns}$ are marked with vertical dashed lines. Right panel: The LFN reduction effect achieved by the median values of parameters. Results shown here are also based on a LFN-only simulation, and the ASD in grey indicates the original readout of science interferometer 12 (detrended). The black dashed lines are the theoretical ASDs of residual LFN calculated with $\Delta R_{\rm median,I/II}$.
@@
In all the aforementioned investigation, for the sake of generality, laser locking was not implemented hence all six lasers are assumed independent of each other. While, in practice, once a specific locking scheme is set, the six lasers become correlated, which would consequently affect the residual LFN after TDI, and the covariances between different arms obtained by TDIR. In Appendix B, we take the locking scheme N1-LA12 [38] as an example to offer a glimpse at the posterior distribution of delay parameters in the presence of laser locking, compared to the “no locking” case. Different correlations between the arms are clearly manifested in FIG. 7. The differences introduced by various locking schemes, as well as the selection of “optimal” scheme in the sense of TDIR remains to be comprehensively explored in future research.

### v4

@@ II.1 TDIR on unsynchronized data
In order to express the same signal in different time frames, for an arbitrary time series $f(\tau)$ (e.g. phase signal, clock reading, delay, etc.) we follow the convention introduced by Ref. [36] to add a superscript to $f$ to denote the time frame of argument $\tau$. For example, $f^{t}(\tau)$ is expressed in TCB, while $f^{\hat{\tau}_{i}}(\tau)$ is timestamped by the clock of S/Ci. TCB is a global time frame for events in the solar system, and the clock times are the only times directly measurable by the clocks (such as USOs) onborad the S/Cs. When describing the same event, the latter differs from the former due to relativistic effects induced by the gravitational field and orbital motions of S/Cs [28] , as well as instrumental imperfections, including long-term, deterministic clock drifts and random jitters. The relationship between clock time and TCB can be expressed as
@@
On the other hand, the “delays” that appear in the expressions of beatnote signals and TDI combinations are defined as the time difference between photon reception and emission, recorded in a certain time frame. The time frame corresponding to LTT (denoted as $d^{t}_{ij}(\tau)$) is TCB. As laser propagates from S/Cj to S/Ci, we have
@@
with $t_{i,{\rm reception}}=\tau$, meaning that the photon arrives at S/Ci when the TCB time is $\tau$. Throughout this paper we follow the prescription of Ref. [37] to calculate LTT to the 1st post-Newtonian order. Meanwhile, the PRNR (dubbed $\hat{d}_{ij}^{\hat{tau}_{i}}(\tau)$, neglecting the technical imperfections such as ranging noise, ambiguity and bias, precisely tracks the onboard clock difference between laser reception and emission:
@@
which incorporates the information of both LTT and differential clock deviations between S/Cs. The first equation is expressed in TCB, and the second gives its transformation to the clock time of receiving S/C, as an application of Eq. (2). To effectively suppress LFN with TDI on unsynchronized data, similar to the “TDI without clock synchronization” scheme [32], the objective of TDIR can be interpreted as searching for the functional form that fits $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ best. However, perfect reconstruction of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$ with a fitting function is not practical due to the presence of random clock jitters in $\delta\hat{\tau}_{i}^{t}$. Functional forms with limited number of parameters (as few as possible to ensure the efficiency of MCMC sampling) can at most fit the slowly-varying, deterministic part of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$. We refer to this variable as the “fiducial” TDIR $\bar{d}_{ij}^{\hat{\tau}_{i}}(\tau)$. To be distinguished, the “parametrized” TDIR that will be fitted using the interferometric data is dubbed $R_{ij}(\tau,\bm{\theta})$. The fiducial TDIR will be treated as the “ground truth” of our searching algorithm, and can be calculated via Eq. (5) assuming that there is no random clock jitters. Meanwhile, the parametrized TDIR contains unknown model parameters to be fitted using the beatnote signals.
@@
Figure 1: A schematic layout of Taiji’s constellation. Each S/C is equipped with two OBs, on which the readouts of science, reference, and test-mass interferometer are generated, including the carriers and sidebands. The indices of signals and noises are the same as the OBs hosting them. S/C2 and S/C3 share the same design of optical measurement system as S/C1. For clarity and conciseness, the specific arrangements of optical components on the OBs are omitted.
The delay operators associated with $d^{t}_{ij}(\tau)$, $\hat{d}^{\hat{\tau}_{i}}_{ij}(\tau)$ and $\bar{d}^{\hat{\tau}_{i}}_{ij}(\tau)$ are denoted as $\textbf{D}_{ij}$, $\textbf{D}_{ij}^{\hat{\tau}_{i}}$and $\bar{\textbf{D}}_{ij}^{\hat{\tau}_{i}}$, respectively. Assuming $f$ is a phase signal generated on S/Cj, after propagated to S/Ci and timestamped by the clock onboard S/Ci as $\tau$, the resulting signal can be described equivalently with the two expressions below [32]:
@@
The left hand side is expressed in TCB, with the arriving TCB time $t=t^{\hat{\tau}_{i}}(\tau)$ determined according to Eq. (2), and the right hand side is expressed in the clock time of S/Ci. The equivalence holds since they describe the same event. On the other hand, according to the definition of $\hat{d}_{ij}^{\hat{\tau}_{i}}(\tau)$, the difference between $\textbf{D}_{ij}^{\hat{\tau}_{i}}f^{\hat{\tau}_{j}}(\tau)$ and $\bar{\textbf{D}}_{ij}^{\hat{\tau}_{i}}f^{\hat{\tau}_{j}}(\tau)$ equals $f$’s time derivative multiplying $\hat{d}^{\hat{\tau}_{i}}_{ij}(\tau)-\bar{d}^{\hat{\tau}_{i}}_{ij}(\tau)$ (i.e. the clock jitter term), which, in the case where $f$ represents some instrumental noise, is a second-order noise term and can be neglected. Therefore, we get a relationship between the delayed noises:
@@
As is explained in Sec. I, for the applications of TDI, it is crucial that the delays used to shift the signals must be consistent with the delays appearing in the equations of signals. Formally Eq. (8-11) looks similar to their “LTT versions” [18, 34, 24, 25], thus all the deduction and algorithm of TDI still applies, and the only two alterations which should be made are to change the delays from LTT to TDIR, and to interpret the beatnote signals as the unsynchronized measurements. We therefore proved that TDIR can be used to suppress LFN on unsynchronized data.
By making slight modifications to the standard formulae in the literature [18, 34, 24, 25], we explicitly present the expressions for the complete TDI data processing flow using TDIR. Firstly, the test-mass-to-test-mass interferometry (denoted as intermediate variable $\xi$) is constructed to eliminate the OB noises:
@@
and the $Y_{2}$ and $Z_{2}$ channels can be obtained by cyclical permutation of the indices : $1\rightarrow 2$, $2\rightarrow 3$, $3\rightarrow 1$. The time frame of each TDI channel will be the clock time of the ending S/C of the virtual optical path (e.g. S/C1 for $X_{2}$), since the conversion of time frames from the emitting S/C to the receiving S/C is also encoded in the TDIR (i.e. the slow-varying part of Eq. (4-5) contains the clock deviations of both S/Cs). At last, the residual clock noise (originating from the $q$ terms in Eq. (8-11)) in the second-generation TDI channels is still 2 - 3 orders larger than the desired noise level, thus an extra step is required to mitigate this noise source. Various clock noise reduction algorithms exist in the literature [34, 35, 39]. Here we represent this step in a unified abstract form:
@@ II.2 Requirements of noise suppression on TDIR
×$\times\left[P_{\rm OMS}(f)+P_{\rm ACC}(f)(\cos 2u+3)\right],$
@@
Next, we investigate the residual LFN, OB noise and clock noise caused by the ranging errors. For generality, we do not set a specific laser locking scheme [42, 38] but rather assume all the six lasers are independent. Besides, to derive analytical expressions for a given level of ranging accuracy, the errors of ranging are set as constant biases dubbed $\Delta R_{ij}$. According to Eq. (12-II.1), the PSDs of the residual LFN and OB noise in the $X_{2}$ channel are
@@
where $\omega\equiv 2\pi f$, $\nu_{0}=281.6\ {\rm THz}$ is the central frequency of laser, and we have defined
@@
as the average ranging error. Eq. (18) is in consistency with the results of previous studies [32, 43], and one can easily verify that a $\mathcal{O}(10)\ {\rm ns}$ requirement is reasonable for the noise models in consideration. While, under the same $\Delta R$, the residual OB noise is much smaller. The algorithms of clock noise reduction varies in the literature (e.g. [34, 35, 39]), thus the residual clock noise can not be written in a unified analytical form. We adopt the treatment of Ref. [35], and explore the impact of ranging error via a clock-noise-only simulation.
The analytical and simulation results are visualized in FIG. 2. In our simulation, the ranging errors are generated according to normal distributions $\mathcal{N}(0,\sigma R)$, with the values of $\sigma R$ vary for different noises, which are 10 ns, 10 ns, 1 ms for the LFN, OB noise and clock noise, respectively. The grey, green and orange amplitude spectral density (ASD) plots represent the simulated residual LFN, OB noise and clock noise, with the corresponding theoretical results shown in black dashed curves. The red solid and dotted lines stand for 1 and 1/10 times the secondary noises, respectively, which are set as the thresholds for TDI noise suppression. Besides the consistency between theoretical analysis and simulations, the figure also indicates that although the ground-tracking ranges deviate from the fiducial TDIR for less than 1 ms, they are already sufficient for suppressing the clock noises. This result is beneficial for the practical implementation of TDIR. Currently the post-TDI clock noise algorithms are usually more complex than TDI itself, therefore require more computational time. To tackle this issue, we do not need to compute $X_{2,q}$ at every iteration of the MCMC run, but rather calculate it with the ground-tracking ranges in advance. Based on this idea, we have designed the whole algorithm of TDIR, which is shown with a simplified flowchart in FIG. 3. Furthermore, within the regions of parameter space of our interest (tens of nanoseconds around the fiducial TDIR), OB and clock noises only make negligible contributions to the total noise budget, thus the performance of TDIR is mainly determined by LFN and the secondary noises.
Figure 3: A simplified flowchart for TDIR. Shown with the light blue parallelogram boxes are the main inputs of the algorithm. The steps in the orange box depend on the undetermined TDIR parameters and should be re-calculated in each iteration of of the MCMC sampling. The dashed arrows represent the clock noise processing steps. As the suppression of clock noise has lower requirements on the accuracy of ranging, the correction term $X_{2,q}$ can be pre-calculated with the ground-tracking ranges, and hence doesn’t need to be included into the loop. This effectively reduces the total computation time of TDIR.
@@ III TDIR in the Bayesian framework
Assuming that the instrumental noises are Gaussian and stationary, $\mathcal{L}$ can be expressed as:
@@
where $\tilde{X}(f;\bm{\theta})$ represents the Fourier transform of an arbitrary TDI channel named $X$, and $P_{X}(f;\bm{\theta})$ is its one-sided PSD (periodogram). The data is whitten by the theoretical one-sided PSD of the secondary noises $P^{\rm 2nd}_{X}(f)$, since secondary noises dominate the TDI data in the region of parameter space of our interest (tens of nanoseconds around fiducial TDIR). In the realistic unequal-arm situation, $P^{\rm 2nd}_{X}(f)$ also depends on the LTTs, which can take the average values of ground-tracking ranges within a day. Using Eq. (41-45), it can be verified that neither the ranging uncertainties ($\sim 0.1\ {\rm ms}$) nor the variations of LTTs during a day ($\sim 1\ {\rm ms}$) have considerable impacts on $P^{\rm 2nd}_{X}(f)$ via numerical or analytical calculation. Shown in the left panel of FIG. 5 are the relative differences of $P^{\rm 2nd}_{X}(f)\ (X\in\{A_{2},E_{2}\})$ caused by the variations of armlengths at 1 ms order. As can be seen from the figure, the accuracy of modeling the secondary noise using the average ground-tracking ranges is typically greater than 99%, and even at the characteristic frequencies ($f=0.025n$ Hz) where the armlengths have the greatest impacts, the accuracy is still above 90%. Eq. (24) is equivalent to the ones adopted by previous studies [24, 25]. In the more general “global fit” [46] scenario where GW signals exist, ${\rm ln}\mathcal{L}$ should be replaced by a more familiar form $-1/2(X-h|X-h)$, $h$ being the templates of GW signals.
Figure 5: Left panel: The relative differences of the theoretical secondary noise PSDs caused by the variations of armlengths at 1 ms order. Right panel: The ASDs of secondary noises and residual LFN in the $A_{2}$ and $T_{2}$ channels, with the armlengths calculated from the orbit data in the 1st day, and the ranging error are set to 10 ns. For both panels the amplitudes of $E_{2}$ are similar to those of $A_{2}$ hence not shown.
Simultaneously estimating the delay parameters and GW source parameters can be a challenging task. Additionally, Taiji will be able to detect the GWs from tens of millions of compact binaries in the Milky Way, and the majority of them cannot be resolved individually, resulting in a foreground confusion noise in the mHz band [47]. Precise modeling of this noise is only possible when all the bright GW signals have been correctly recognized and subtracted. Therefore, to reduce the encounters with GW signals and foreground noise, we set $f_{\rm min}=10\ {\rm mHz}$, above the most sensitive band of Taiji. This choice of $f_{\rm min}$ also allows us to accelerate the estimation process by incorporating a preliminary search phase, see Sec. IV. Ideally, under the condition that the computing capacity is sufficient or the algorithm has been optimized in speed, one may discard this phase and raise $f_{\rm min}$ above 0.1 Hz to further minimize the impact of GWs.
@@
and the total logarithmic likelihood function is the sum of $\{A_{2},E_{2},T_{2}\}$ channels. The PSDs of secondary noises in the optimal channels can be found in Appendix A, where we have also presented the PSDs of residual LFN due to the coupling with ranging errors. Observing Eq. (25), the contribution of a certain TDI channel to the total likelihood is determined by the ratio between residual LFN and secondary noises. These two terms for the $\{A_{2},T_{2}\}$ channels are shown in the right panel of FIG. 5, with the armlengths calculated from the numerical orbit data in the 1st day, and the ranging errors are set to 10 ns. The amplitudes of noises in $E_{2}$ are similar to those of $A_{2}$ hence not shown. As can be seen from the figure, within the frequency band in concern, each channel contributes similarly to the likelihood.
We know from the explanation in Sec. II that the timestamps of each Michelson channel is determined by the clock at the end of the virtual optical path. Some comments should be made about the combination of these “roughly synchronized” channels. Firstly, taking the $A_{2}$ channel as an example and assuming the desynchronization between S/C1 and S/C3 is a constant $\delta t$, then the error in modeling $P_{A_{2}}^{\rm 2nd}$ with Eq. (43) is approximately $2\omega\delta t{\rm Im}(Z_{2}X_{2}^{*})$, much smaller than the other terms of $P_{A_{2}}^{\rm 2nd}$ under the condition of $\omega\delta t\ll 1$, which is met across the whole 0.1 mHz - 1 Hz band. Secondly, the “best fit” delays resulting from different TDI channels could mismatch due to the desynchronization of corresponding S/Cs. Ref. [51] summarized the criteria for S/C movements to be qualified for space-based GW detection: the relative velocities between S/Cs should be smaller than 5 m/s for LISA and 6 m/s for Taiji. we take a conservative value of 10 m/s, which yields the variations of delays smaller than a nanosecond for $\delta t<0.03\ {\rm s}$. Therefore, we conclude that a 0.1 ms order desynchronization between different Michelson channels does not impact the combination into optimal channels.
@@ IV.2 Results of MCMC sampling
Figure 7: Left panel: The distributions of $\Delta R$ calculated from the MCMC samples of Simulation I, II. $\Delta R_{\rm median,I}=0.49\ {\rm ns}$ and $\Delta R_{\rm median,II}=5.28\ {\rm ns}$ are marked with vertical dashed lines. Right panel: The LFN reduction effect achieved by the median values of parameters. Results shown here are also based on a LFN-only simulation, and the ASD in grey indicates the original readout of science interferometer 12. The black dashed lines are the theoretical ASDs of residual LFN calculated with $\Delta R_{\rm median,I/II}$.
@@
In all the aforementioned investigation, for the sake of generality, we have assumed that laser locking is not implemented hence all six lasers are independent of each other. While, in practice, once a specific locking scheme is set, the six lasers become correlated, which would consequently affect the residual LFN after TDI, and the covariances between different arms obtained by TDIR. Taking the locking scheme N1-LA12 [38] as an example, for the equal-arm case, the residual LFN in the Michelson $X_{2}$ channel caused by ranging errors reads:
$P^{\rm L}_{X_{2}}(f;\Delta R_{ij})$ $=$ $16\sin^{2}(2u)\sin^{2}(u)\omega^{2}P_{\rm L}(f)$ (35)
$\times$ $\left(\Delta R_{13}+\Delta R_{31}-\Delta R_{12}-\Delta R_{21}\right)^{2}.$
And the PSD of the secondary noises remains the same as Eq. (17). With Eq. (27), one can easily verify that the Fisher information matrix is singular, indicating that there is degeneracy among the parameters. In Appendix B, we offer a glimpse at the posterior distributions of delay parameters in the “locked lasers” case, compared to the “no locking” case. The figure shows that the combination of different TDI channels can mitigate the degeneracy to a certain extent, but the correlation between different arms still exists. Furthermore, the differences introduced by various locking schemes, as well as the selection of “optimal” scheme in the sense of TDIR remains to be comprehensively explored in future research.